ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53A05, 53A10 Hertrich-Jeromin, Udo, TU Berlin (Germany) Transformations of Isothermic surfaces Certain transformations for isothermic surfaces will be presented, as well as their interrelations. Relations with the integrable system approach to isothermic surfaces, with discrete isothermic nets, and with classical transformations of minimal and cmc surfaces in spaces of constant curvature will be discussed.